Find all p which satisfy both the inequalities 0>=54p-144 and 0>12-20p. Express your answer in interval notation, reducing any fractions in your answer.
Answer by macston(5194)

.
0>=54p-144
144>=54p
144/54>=p
8/3>=p
.
0>12-20p
20p>12
p>12/20
p>3/5
.
3/5 < p < =8/3
.
(3/5,8/3]